OpenStudy (darkbeautystar):

In physics, we use the equation s= d/t to show the speed of an object. This equations means that the speed of an object is equal to distance it moved divided by the time it moved. If a car was going 65 miles per hour and went 910 miles; how long was it driving?

OpenStudy (darkbeautystar):

@phi

OpenStudy (darkbeautystar):

one more q after this

OpenStudy (darkbeautystar):

oh man--there is one more question for the 1st one we were working on

OpenStudy (phi):

we can go back to that other post.

OpenStudy (darkbeautystar):

i closed that post

OpenStudy (phi):

For this problem, you have s= d/t and you are told speed is 65 and distance is 910 where do we put the 65 in s= d/t i.e. which letter stands for speed?

OpenStudy (darkbeautystar):

if you just help me with these 2 last questions we can figure something out

OpenStudy (darkbeautystar):

s

OpenStudy (phi):

ok, so replace the s with 65 what does the formula look like ?

OpenStudy (darkbeautystar):

65=d/t

OpenStudy (darkbeautystar):

what grade are u in anyway?

OpenStudy (phi):

yes. you also know the distance is 910 which letter stands for distance? where should the 910 go?

OpenStudy (darkbeautystar):

d--so 65=910/t

OpenStudy (phi):

yes. now you have to find what number t is

OpenStudy (darkbeautystar):

yep

OpenStudy (phi):

multiply both sides by t that means write *t on both sides can you do that ?

OpenStudy (darkbeautystar):

hangon-- i might be able to get his answer- give me a sec to figure this out

OpenStudy (darkbeautystar):

t = 14 ?

OpenStudy (phi):

you should try now while you have help

OpenStudy (darkbeautystar):

i used a calc to divide

OpenStudy (phi):

yes, but how did you know you should divide ?

OpenStudy (darkbeautystar):

i ment * sorry-- my auto correct has been putting in words if i use them oftenly

OpenStudy (darkbeautystar):

* = multiply

OpenStudy (darkbeautystar):

i have just learned from you and my school

OpenStudy (phi):

you solve by doing this 65= 910/t multiply both sides by t 65t= 910 * t/t on the right side t divided by itself is 1 (that is why we multiplied by t) 65 t= 910 now divide both sides by 65 you hget 65/65 * t = 910/65

OpenStudy (phi):

that simplifies to t= 910/65 t= 14

OpenStudy (darkbeautystar):

yep-- thats what i said ;)

OpenStudy (phi):

the point is not just to get the answer, it is to know how to get the answer.

OpenStudy (darkbeautystar):

but how would i right that out for the answer

OpenStudy (darkbeautystar):

well i know

OpenStudy (darkbeautystar):

@phi

OpenStudy (phi):

you would write out the lines I posted.

OpenStudy (darkbeautystar):

ok

OpenStudy (darkbeautystar):

@phi Misha wants to know how many pepperonis to put on his pizza. He was told that he should use 1 pepperoni rounded to the nearest whole pepperoni for each .75 inches of diameter in a pizza. Show an equation that would solve how many pepperonis he would need if his pizza has a 8 inch radius.

OpenStudy (phi):

you need a formula for the area of a circle do you know it?

OpenStudy (darkbeautystar):

no

OpenStudy (phi):

can you look in your notes?

OpenStudy (darkbeautystar):

ohh i found it i do know how

OpenStudy (darkbeautystar):

my teacher taught me a way to do it. so i need the numbers and i can calculate in my notes

OpenStudy (darkbeautystar):

it is really cool

OpenStudy (darkbeautystar):

never mind